# Masked Image Modeling via Dynamic Token Morphing

### v1

@@ Appendix C Ablation study on loss functions
Appendix D Implementation details
Pre-training on ImageNet-1K.
The pre-training recipe for DTM mainly follows the recipe of BEiT v2 [28]. Table 12 reports the implementation details for pre-training. We adopt the hyperparameters for ViT-B/16 and ViT-L/16 pre-training from BEiT v2. Specifically, we use layer scales of 0.1 and for ViT-B/16 and ViT-L/16, respectively. We employ both relative positional embeddings and shared relative positional embeddings. The maximum gradient value is constrained to 3.0. We apply color jittering followed by random resizing and cropping for data augmentation. The hyperparameters for ViT-S/16 replicate the settings of ViT-B/16.
Fine-tuning on ImageNet-1K.
The fine-tuning recipe for DTM mainly follows the standard supervised learning recipe for ViT. Specifically, we adopt commonly used values for RandAugment, Mixup, Cutmix, and Label Smoothing On the other hand, we employ relative positional embeddings. Stochastic depth is applied with values of 0.1, 0.1, and 0.2 for ViT-S/16, ViT-B/16, and ViT-L/16, respectively The overall recipe is detailed in Table 13.
Fine-tuning on ADE20K.
Table 14 summarizes the fine-tuning recipe of ViT/16 for the semantic segmentation task on ADE20K [50]. We employ AdamW with momentum (0.9, 0.999) and warm-up for 1500 iterations. The learning rate is linearly scheduled with a value of . We apply layer-wise learning rate decay of 0.75, stochastic depth of 0.1, and weight decay of 0.05. The model is fine-tuned using 8 V100-32GB GPUs.

### v2

@@ Appendix 0.B Compatibility with Superpixel Algorithms
0.B.1 Ablation Study on the Number of Schedules
Our method can further enhance the diversity of target morphed tokens by employing multiple schedules within a single iteration. Thus, we study the effects of learning diverse morphed tokens derived from multiple morphing schedules simultaneously. We compare the fine-tuning accuracy of models pre-trained by our DTM, varying the number of schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [39] and ImageNet-100 [39]. Table B reveals that exploring diverse target morphed tokens improves the representation capability of pre-trained models, leading to increased fine-tuning accuracies of at least 0.1p and 0.2p on ImageNet-1K and ImageNet-100 compared to the model pre-trained by the single DTM approach, respectively. However, the goal of experiencing diverse morphed tokens at once appears to be attained with double scheduling, resulting in no additional gains in performance through further exploration. Given that utilizing two schedules yields the best and is most efficient among all other multiple scheduling options, we adopt the double morphing scheduling approach.
Appendix 0.C Additional ablation studies
0.C.1 Ablation Study on Randomness of DTM.
0.C.1.1 Randomness in the number of morphing tokens.
Our DTM randomizes the number of morphing tokens since fixing the number of morphing tokens does not adequately generate diverse morphed tokens. To verify the effectiveness of the randomness, we compare the fine-tuning accuracies of models pre-trained using DTM with random and fixed numbers of morphing tokens. We pre-train and fine-tune the models for 100 epochs on ImageNet-100 [39]. As reported in the 2nd and 5th rows of Table C, exploring diverse morphed tokens improves the fine-tuning accuracy by 0.5p. This result demonstrates the impact of varying numbers of morphing tokens.
0.C.1.2 Randomness in gradual token morphing.
As morphed tokens can vary by the number of morphing iterations, we apply randomness in token morphing iterations. We compare fine-tuning accuracies of the pre-trained models with random and fixed iteration numbers for morphing to validate the effect of randomness. As shown in the 3rd and 5th rows of Table C, randomly selecting the iteration number for morphing enhances the performance, confirming its effectiveness.
0.C.2 Ablation Study on target normalization.
Target normalization is proven to have a significant impact on MAE [19]. Thus, we verify the impact of target normalization on our DTM. We employ ViT-B/16 with a resolution of for comparison. As shown in the 4th and 5th rows of Table C, target normalization does not yield a positive effect on our DTM, resulting in a decrease in fine-tuning accuracy from 87.9 to 87.7.
Appendix 0.D Implementation details
0.D.0.1 Pre-training on ImageNet-1K.
The pre-training recipe for DTM mainly follows the recipe of BEiT v2 [34]. Table D reports the implementation details for pre-training. We adopt the hyperparameters for ViT-B/16 and ViT-L/16 pre-training from BEiT v2. Specifically, we use layer scales of 0.1 and for ViT-B/16 and ViT-L/16, respectively. We employ both relative positional embeddings and shared relative positional embeddings. The maximum gradient value is constrained to 3.0. We apply color jittering followed by random resizing and cropping for data augmentation. The hyperparameters for ViT-S/16 replicate the settings of ViT-B/16.
0.D.0.2 Fine-tuning on ImageNet-1K.
The fine-tuning recipe for DTM mainly follows the standard supervised learning recipe for ViT. Specifically, we adopt commonly used values for RandAugment, Mixup, Cutmix, and Label Smoothing. On the other hand, we employ relative positional embeddings. Stochastic depth is applied with values of 0.1, 0.1, and 0.2 for ViT-S/16, ViT-B/16, and ViT-L/16, respectively. The overall recipe is detailed in Table E.
@@ Appendix 0.D Implementation details
0.D.0.3 Fine-tuning on ADE20K.
Table F summarizes the fine-tuning recipe of ViT/16 for the semantic segmentation task on ADE20K [55]. We employ AdamW with momentum (0.9, 0.999) and warm-up for 1500 iterations. The learning rate is linearly scheduled with a value of We apply layer-wise learning rate decay of 0.75, stochastic depth of 0.1, and weight decay of 0.05. The model is fine-tuned using 8 V100-32GB GPUs.
0.D.0.4 Transfer learning.
We follow the fine-tuning recipes for DTM to conduct transfer learning to iNaturalist datasets, including iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 [45] and FGVC datasets, including Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs. However, we additionally perform grid searches of learning rates and weight decay. Specifically, we fine-tune the models with learning rates of , , and and weight decays of 0.05 and 0.1.